1. If we=1,000 represents a two-period consumer's lifetime wealth and r=0.05 denotes the real rate of interest, the slope of the consumer's budget line is equal to

Question 12 options:

a. -1.05

b. 50.05

c. -0.95

d. -0.00005

e. 50

2.

Suppose that the utility function of the consumer is U(c,c’)=ln(c) + b ln( c’ ) . Analyze the solution of the model to show show that optimal consumption increases over time when

Question 3 options:

a. b > 1

b. b (1+r)<1

c. b (1+r)<0

d. b (1+r)=1

e. b (1+r)>1
Answer #1

Q1) option a)

the Slope of intertemporal budget constraint = -(1+r)

= -1.05

Q2) option e)

at eqm in intertemporal utility maximization problem

MRS = (1+r)

MRS = MUc/MUc`

= C`/bC

So , C`/bC = (1+r)

C`/C = b(1+r)

So Consumption will increase over time if

C`>C

C`/C > 1

So, b(1+r) > 1